# Confidence Calibration in Vision-Language-Action Models

### v1

@@ 4.5 Calibration Across Action Dimensions
Token-based VLAs such as OpenVLA or RT-2-X decompose a low-level command into 7 different action dimensions, i.e. Cartesian displacements, yaw-pitch-roll rotations, and a binary gripper toggle, predicted sequentially by an image-conditioned autoregressive LLM. Our baseline confidence estimate collapses this structure into a single scalar by averaging the top-token probabilities across the dimensions (Equation (3)); implicitly, this assumes that every dimension is calibrated to the same degree. That assumption may not hold in practice: a gripper ‘open’ or ‘close’ token may appear in nearly every demonstration, whereas a $90^{\circ}$ wrist roll could be rare. Such dataset imbalances and other implementation details might skew calibration across dimensions, and our relevant algorithms, e.g., for post hoc recalibration, might benefit from explicitly addressing this.
To study this question, our experiment consists of two parts:
  1. 1.
    Per-dimension calibration audit. For each dimension $d$ we treat $\max_{k}p^{(d)}_{t,k}$ as that dimension’s confidence and compute $\text{ECE}_{1}$, to examine whether different degrees of freedom are differentially calibrated.
  2. 2.
    Targeted recalibration. We compare classic Platt scaling to the action-wise variant of Section 3.6, which learns independent scaling parameters for each action dimension before averaging the transformed confidences across dimensions.
@@
We perform our study using the baseline confidence estimates and all fine-tuned model variants (full, Quant-8, Quant-4) on the Spatial and Goal task suites. Each experiment is run for 1000 trials, with random 20%/80% calibration-test splits. Given the reduced size of the test set, we measure calibration with 10 equal-mass bins (instead of 12 as in other experiments). Beyond traditional Platt scaling, we also include a temperature scaling (Guo et al., 2017) baseline, where a single temperature is applied to all dimensions’ logits before the softmax. Figure 9 shows results with the baseline estimates for the full precision OpenVLA models; Appendix Figure 19 includes the results with the quantized models.
Two key observations arise:
  1. 1.
    $\text{ECE}_{1}$ varies by up to 5 times across dimensions, with no consistent best or worst dimensions across settings. It follows that a single scalar confidence masks significant differences in dimension-wise confidence estimates.
  2. 2.
    Replacing global Platt scaling with per-action-dimension transforms consistently lowers calibration error (and also improves over the temperature scaling baseline). ECE is improved by over 20% in some cases, without changing the selected tokens or adding meaningful runtime overhead.
@@
Together, these results indicate that VLA calibration research should treat each degree of freedom as its own concern in order to improve effectiveness of relevant algorithmic tools. Dimension-aware post hoc methods such as action-wise Platt scaling offer a simple yet effective path toward that goal. More broadly, our findings underscore that calibrating VLAs introduces challenges not seen in other domains, so meaningful progress will demand substantial domain-specific research rather than wholesale adoption of understanding and techniques from other areas of deep learning.
@@
Figure 9: On the left of each pair of plots, we compare miscalibration across action dimensions. On the right, we compare the performance of typical Platt scaling (and temperature scaling) to action-wise Platt scaling.
@@ 5 Discussion
Our work set out to answer a simple but previously unexplored question: can existing VLA policies tell us, in a statistically reliable way, how likely they are to succeed? Our experiments offer some preliminary answers. First, we find that better task performance correlates with lower calibration error, indicating that optimization for task performance and reliable confidence estimation are not in tension. Second, a zero-training algorithm using an ensemble of rephrased natural language instructions reduces expected calibration error by similar-to\!$\sim$20–40across 3 LIBERO suites and OpenVLA variants. This suggests that some miscalibration is simply high-variance noise in the language channel, and highlights a direction for future work in more reliable action prediction via perceptual perturbation. Third, our results indicate that miscalibration is not uniform across potential sources of confidence estimates. Instead, it varies over time and across action dimensions, and adaptive post hoc fixes (e.g., time-horizon-aware monitoring and action-wise Platt scaling) can exploit that structure. Together, our findings help establish a research agenda for VLA calibration: understanding how task accuracy and calibration co-evolve across diverse environments, leveraging natural perceptual variance via ensembles, incorporating temporal and per-dimension confidence dynamics, and developing rigorous algorithms that exploit these structures for identifying and addressing uncertainty.
@@ 6 Limitations
All experiments use OpenVLA fine-tuned in the LIBERO simulator. While this choice enables the controlled resets and hundreds of rollouts per task suite needed for stable calibration estimates, it limits architectural and domain diversity. However, running the same study on physical robots would require thousands of real executions, which can be prohibitively time-consuming and expensive. Simulation therefore may offer the most practical testbed for calibration research for the foreseeable future, although it cannot accurately model the sensor noise, latency, wear-and-tear, or other physical factors that may influence confidence calibration in the real world.
@@ 7 Future Work
Several extensions follow naturally from our work. Confidence calibration should be benchmarked across other environments, robot embodiments, and VLA architectures, including diffusion-based planners, continuous regression heads, and flow-matching controllers. Establishing strong baselines for these architectures will likely require new confidence surrogates and post hoc adjustments. Likewise, perturbation techniques besides instruction paraphrasing, e.g., inserting synthetic lighting changes or random distractors, could expose additional failure modes and inspire new multimodal ensemble techniques. Running such experiments on real robots will be essential for understanding how various physical factors affect VLA calibration in the real world. Finally, our temporal analysis hints at a period mid-trajectory where confidence is most trustworthy; integrating that signal into selective execution, planning, or human-in-the-loop systems is an open challenge for designing adaptive risk-aware robotic pipelines.
Overall, we hope that our work lays the groundwork for the rich body of research on calibration in VLAs necessary to place these models at the center of trustworthy real-world robotics systems.
@@ Appendix A Additional Experiment Details
A.1 Rephrasing Prompts
The prompts given to GPT-4o-mini for rephrasing instructions from the LIBERO robot simulation environment are listed in Tables 5 and  6. Prompt 1 is the main prompt used throughout the experiments, and Prompts 2 and 3 are used to ablate the sensitivity to the rephrasings. Note that the different prompts lead to substantially different rephrasings, as Prompts 2 and 3 mandate the retention of the words “pick” and “place”, while as Table 1 shows, Prompt 1 leads to these words being replaced.
@@ B.1 Relationship Between Task Success and Calibration
Figure 10 shows the relationship between calibration error and task error using the prompt ensemble method for estimating confidence.
Figure 10: Comparison of (4 different) calibration error measurements against task error rates for all model variants and task suites, using the Reprompt method to produce confidence estimates. A clear trend emerges: lower task error rates are associated with lower calibration error.
@@ B.2 Prompt Ensembling
Figure 11 shows trends in percent change in calibration error using the prompt ensemble method. Figure 12 offers results for the ablation across ensemble sizes using the Goal and Object task suites.
@@
Quant-4 Spatial Baseline 0.067 0.086 0.149 0.503
@@
Table 7: Calibration error measurements using the 4-bit fine-tuned model version and 3 different LIBERO task suites, for 2 different methods of confidence estimation: (1) baseline (2) ensembling over semantically equivalent prompts (“Reprompt”). The prompt ensemble method improves all measures across all settings.
@@
Figure 12: Ablating the number of prompts in the prompt ensemble for the Goal and Object task suites. Increasing the number of prompts generally improves calibration error.
B.3 Calibration Across Task Time
Figures 13, 14, 15, 16, and 17 offer additional results on calibration across the task time horizon. Figure 18 shows further qualitative examples from the context-aware monitoring experiment.
@@ B.3 Calibration Across Task Time
B.4 Calibration Across Action Dimensions
@@
Figure 19: On the left of each pair of plots, we compare miscalibration across action dimensions. On the right, we compare the performance of typical Platt scaling (and temperature scaling) to action-wise Platt scaling.

### v2

@@ 4.5 Calibration Across Action Dimensions
Some token-based VLAs such as OpenVLA or MolmoAct decompose a low-level command into tokens explicitly corresponding to their 7 different action dimensions. Our baseline confidence estimate collapses this structure into a single scalar by averaging the top-token probabilities across the dimensions; implicitly, this assumes that every dimension is calibrated to the same degree. That assumption may not hold in practice: a gripper ‘open’ or ‘close’ token may appear in nearly every demonstration, whereas a $90^{\circ}$ wrist roll could be rare. Such dataset imbalances and other implementation details might skew calibration across dimensions, and relevant algorithms, e.g., for post hoc recalibration, might benefit from explicitly addressing this. To study this question, our final experiment consists of two parts. (1) Per-dimension calibration audit. For each dimension $d$ we treat $\max_{k}p^{(d)}_{t,k}$ as that dimension’s confidence and compute $\text{ECE}_{1}$, to examine whether different degrees of freedom are differentially calibrated. (2) Targeted recalibration test. We compare classic Platt scaling to the action-wise variant of Section 3.6, which learns independent scaling parameters for each action dimension before averaging the transformed confidences across dimensions.
@@
We perform our study using MolmoAct and all 3 OpenVLA variants (full, Quant-8, Quant-4) fine-tuned on the Spatial and Goal task suites. Each experiment is run for 1000 trials, with random 20%/80% calibration-test splits. Given the reduced size of the test set, we measure calibration with 10 equal-mass bins (instead of 12 as in other experiments). Beyond traditional Platt scaling, we also include a range of baselines: temperature scaling (guo2017calibration), histogram binning (zadrozny2001obtaining), and Platt binning (kumar2020verifieduncertaintycalibration). (Note: temperature scaling is excluded from MolmoAct results, as the variable-length action encodings preclude a straightforward adaptation.) Figure 6 shows results with the baseline estimates for the full precision OpenVLA and MolmoAct models; Appendix Figure 19 includes the results with the quantized models.
@@
Two key observations arise. First, $\text{ECE}_{1}$ varies by up to 2 times across dimensions, with no consistent best or worst dimensions across settings. It follows that a single scalar confidence masks significant differences in dimension-wise confidence estimates. Second, we can observe that replacing global Platt scaling with per-action-dimension transforms consistently lowers calibration error (and also improves over the additional baselines). ECE is improved by over 20% in some cases, without changing the selected tokens or adding meaningful runtime overhead. Together, these results indicate that VLA calibration research should treat each degree of freedom as its own concern in order to improve the effectiveness of relevant algorithmic tools. Dimension-aware post hoc methods such as action-wise Platt scaling offer a simple yet effective path toward that goal. More broadly, our findings underscore that calibrating VLAs introduces challenges not seen in other domains, so meaningful progress will demand substantial domain-specific research rather than wholesale adoption of understanding and techniques from other areas of deep learning.
@@
Figure 6: On the left of each pair of plots, we compare miscalibration across action dimensions. On the right, we compare the performance of baselines to action-wise Platt scaling.
@@ 5 Discussion
In this work, we set out to answer a simple but previously unexplored question: can existing VLA models tell us, in a statistically reliable way, how likely they are to succeed? Our experiments offer some preliminary answers. First, we find that average top-token probability is a reasonable confidence baseline, and that decisions like architecture and learning objectives may affect the crucial relationship between task performance and calibration error. Second, a zero-training algorithm using an ensemble of rephrased natural language instructions reliably reduces expected calibration error across 3 LIBERO suites and 3 VLA variants. This suggests that some miscalibration is simply high-variance noise in the language channel, and highlights a direction for future work in more reliable action prediction via input perturbation. Third, our results indicate that miscalibration is not uniform across potential sources of confidence estimates. Instead, it varies over time and across action dimensions, and adaptive post hoc fixes (e.g., time-horizon-aware monitoring and action-wise Platt scaling) can exploit that structure. Together, our findings point towards some interesting initial directions of study for VLA calibration: understanding how task accuracy and calibration co-evolve across diverse modeling approaches and environments, creating calibration tools grounded in practical VLA implementation details, incorporating temporal confidence dynamics, and developing rigorous algorithms that exploit these structures for identifying and addressing uncertainty.
@@ 6 Limitations
All experiments in this work are conducted in simulation. This choice is common in robotics research (e.g., lillicrap2015continuous; schulman2017proximal), and enables the controlled resets and hundreds of rollouts per task suite needed for stable calibration estimates. However, it limits domain diversity and leaves open questions of how these results translate to real settings. Replicating the same study on physical robot hardware would require thousands of physical executions, which can be prohibitively time-consuming and expensive. Accordingly, simulation remains the most practical testbed for calibration research in the near term, though it cannot fully capture sensor noise, latency, wear-and-tear, and other physical factors that may influence confidence calibration in the real world. Studying these questions and techniques on real robots will ultimately be essential for understanding how these and other factors affect VLA calibration under realistic conditions.
@@ 7 Future Work
Several extensions follow naturally from our work. Confidence calibration should be benchmarked across other environments, robot embodiments, and VLA architectures, including diffusion-based planners, continuous regression heads, and flow-matching controllers. Establishing strong baselines for these architectures will likely require new confidence surrogates and post hoc adjustments. Likewise, perturbation techniques besides instruction paraphrasing, e.g., inserting synthetic lighting changes or random visual distractors, could expose additional failure modes and inspire new multimodal ensemble techniques. Running such experiments on real robots will be essential for understanding how various physical factors affect VLA calibration in the real world. Uncertainty estimation also holds the potential to guide efficient data gathering and fine-tuning, as in active learning (Wang_2017). Further, while all of our experiments focused on fine-tuned in-distribution settings, future work should consider these phenomena in the zero-shot and out-of-distribution settings. Finally, our temporal analysis hints at a period mid-trajectory where confidence is most trustworthy; integrating that signal into selective execution, planning, or human-in-the-loop systems is an open challenge for designing adaptive risk-aware robotic pipelines.
Overall, we hope our work helps build a foundation for the research on VLA calibration needed to make these models trustworthy components of real-world robotic systems.
@@ Appendix A Additional Experiment Details
Task examples from the LIBERO task suites are shown in Table 3.
Dataset Example Image Example Instruction
Goal [Uncaptioned image] Open the middle drawer of the cabinet.
10 [Uncaptioned image] Put both the alphabet soup and the tomato sauce in the basket.
Table 3: Task examples from the LIBERO task suites: Spatial, Object, Goal, and 10.
A.1 Confidence Estimation with VLA Variants
MolmoAct
At each timestep, MolmoAct predicts a sequence of discrete action tokens for every action dimension, with a softmax distribution over the 256 action bins for each token. For a given action dimension, we take the probability assigned to the executed bin for each of its tokens and average these probabilities to obtain a per-dimension confidence. This yields one scalar confidence value per action dimension. We then average these per-dimension confidences across all action dimensions to obtain a single scalar confidence for the action at that timestep.
UniVLA
At each timestep, UniVLA predicts a fixed number (4) of discrete latent action tokens, each chosen from a codebook of size 16. For each latent token, we take the softmax probability assigned to the selected code as its token-level confidence. Because these latent tokens are jointly decoded into all continuous control dimensions, the confidence scores do not correspond to particular action dimensions. Finally, we average these per-token confidences to obtain a single scalar confidence for UniVLA at that timestep.
NORA
At each timestep, NORA predicts an action chunk encoded as a sequence of discrete tokens, each selected via a softmax over the augmented action vocabulary. For every token, we take the softmax probability of the sampled token as its token-level confidence. Because the NORA tokenizer mixes information from all control dimensions, we do not derive per-dimension confidences. Instead, we compute a single scalar confidence for NORA by averaging these token-level confidences across all tokens in the predicted action chunk for that timestep.
A.2 Rephrasing Prompts
The prompts given to GPT-4o-mini for rephrasing instructions from the LIBERO robot simulation environment are listed in Tables 4 and 5. Prompt 1 is the main prompt used throughout the experiments, and Prompts 2 and 3 are used to ablate the sensitivity to the rephrasings. Note that the different prompts lead to substantially different rephrasings, as Prompts 2 and 3 mandate the retention of the words “pick” and “place”, while as Table 1 shows, Prompt 1 leads to these words being replaced.
@@ Appendix B Additional Experiment Results
All task success and calibration results for 22 VLA/task suite combinations are shown in Table 6.
@@ B.1 Relationship Between Task Success and Calibration
Table 6 features all results plotted in Figure 3.
Dataset Model Quant $\text{ECE}_{1}$ $\text{ECE}_{2}$ Brier NLL Succ. Rate
Spatial NORA - 0.099 0.115 0.105 0.386 0.894
Object NORA - 0.210 0.229 0.090 0.343 0.960
Goal NORA - 0.065 0.077 0.111 0.391 0.880
10 NORA - 0.184 0.212 0.260 0.811 0.674
Spatial MolmoAct - 0.113 0.123 0.134 0.625 0.860
Object MolmoAct - 0.046 0.059 0.062 0.296 0.936
Goal MolmoAct - 0.129 0.141 0.146 0.657 0.846
10 MolmoAct - 0.175 0.184 0.205 0.716 0.772
Spatial UniVLA - 0.162 0.193 0.064 0.282 0.972
Object UniVLA - 0.082 0.135 0.054 0.224 0.962
Goal UniVLA - 0.116 0.163 0.066 0.261 0.958
10 UniVLA - 0.134 0.174 0.092 0.339 0.932
Spatial OpenVLA - 0.088 0.106 0.150 0.533 0.828
Object OpenVLA - 0.060 0.073 0.108 0.401 0.880
Goal OpenVLA - 0.151 0.170 0.207 0.707 0.758
10 OpenVLA - 0.381 0.390 0.398 1.382 0.532
Spatial OpenVLA Quant-8 0.070 0.093 0.138 0.488 0.844
Object OpenVLA Quant-8 0.057 0.065 0.110 0.409 0.876
Goal OpenVLA Quant-8 0.140 0.163 0.205 0.713 0.760
Spatial OpenVLA Quant-4 0.067 0.086 0.149 0.503 0.822
Object OpenVLA Quant-4 0.091 0.112 0.135 0.489 0.850
Goal OpenVLA Quant-4 0.161 0.182 0.231 0.748 0.712
Table 6: All task success and calibration results for 22 VLA/task suite combinations.
@@ B.2 Prompt Ensembling
Table 7 has prompt ensemble results for the quantized OpenVLA models. Figure 7 shows trends in percent change in calibration error using the prompt ensemble method.
@@
OpenVLA Quant-8 Spatial Baseline 0.070 0.093 0.138 0.488
Reprompt 0.050 0.062 0.131 0.434
Object Baseline 0.057 0.065 0.110 0.409
Reprompt 0.041 0.052 0.108 0.375
Goal Baseline 0.140 0.163 0.205 0.713
Reprompt 0.117 0.152 0.194 0.608
OpenVLA Quant-4 Spatial Baseline 0.067 0.086 0.149 0.503
@@
Table 7: Calibration error measurements using the 8-bit and 4-bit fine-tuned model versions and 3 different LIBERO task suites, for 2 different methods of confidence estimation: (1) baseline (2) ensembling over semantically equivalent prompts (“Reprompt”). The prompt ensemble method improves all measures across all settings.
@@
Figure 9: Ablating the number of prompts in the prompt ensemble for the Spatial, Goal, and Object task suites. Increasing the number of prompts generally improves calibration error.
B.3 Ablations for Prompt Ensembles
To understand the robustness of our prompt ensemble results, we perform multiple ablations using the OpenVLA model.
First, we consider the effect of changing the prompt given to GPT-4o-mini for producing the 20 instruction rephrasings. In addition to the prompt used to produce the results in Table 2, we try two additional prompts, recorded in Appendix Table 4 and Table 5; “Prompt 1” is the original prompt, and “Prompt 2” and “Prompt 3” represent new prompts for the ablation. Note that the different prompts lead to substantially different rephrasings, as Prompts 2 and 3 mandate the retention of the words “pick” and “place”, while as Table 1 shows, Prompt 1 leads to these words being replaced. The ablation is run across all 3 task suites using fine-tuned OpenVLA, and we measure $\text{ECE}_{1}$ and NLL. Results are shown in Figure 8. Improvements in calibration error from the Reprompt method are robust to different rephrasing prompts, with all ensemble variants producing lower error than the baseline method across both metrics and all 3 task suites.
Second, we consider the effect of the number of prompts used in the prompt ensemble. Given a sound ensembling approach, we would expect to see improvement as more prompts are added to the ensemble (up to some point). To study this question, we record results using $k\in\{1,5,10,20\}$ different rephrasings (randomly chosen for 1000 trials when $k<20$). Results for the Spatial, Goal, and Object task suites are shown in Figure 9. We see that the algorithm behaves favorably, where calibration error generally decreases as more instructions are included in the ensemble.
B.4 Calibration Across Task Time
Figures 10, 11, 12, 13, and 14 offer additional results on calibration across the task time horizon using the OpenVLA model. Figures 15, 16, and 17 offer results for the same experiment on UniVLA. Figure 18 shows further qualitative examples from the context-aware monitoring experiment.
@@ B.4 Calibration Across Task Time
Figure 15: Empirical study of calibration error across task time horizon for the Spatial task suite and the UniVLA model.
Figure 16: Empirical study of calibration error across task time horizon for the Object task suite and the UniVLA model.
Figure 17: Empirical study of calibration error across task time horizon for the Goal task suite and the UniVLA model.
B.4.1 Qualitative Examples of Context-Aware Confidence Monitoring
Additional qualitative examples of our confidence-aware monitoring demonstration are provided in Figure 18. These include cases where the strategy succeeds and others where it fails (e.g., confidence falls below the threshold while grasping in an ultimately successful trial).
@@ B.4.1 Qualitative Examples of Context-Aware Confidence Monitoring
B.5 Calibration Across Action Dimensions
@@
Figure 19: On the left of each pair of plots, we compare miscalibration across action dimensions. On the right, we compare the performance of typical Platt scaling, temperature scaling, histogram binning, and Platt binning to action-wise Platt scaling.